OpenStudy (anonymous):

Which equation models this problem? Chelsea paid $54.90 for nine bath towels. What is the unit price, p, per towel?

OpenStudy (anonymous):

p/9 = 54.90 9p = 54.90 54.90p = 9 p = 9 ------ 54.90

sammixboo (sammixboo):

Because is Chelsea paid $54.90 for nine bath towels. All that you need to do is take 9 and divide it by 54.90.

sammixboo (sammixboo):

What do you think it is?

OpenStudy (anonymous):

um idek its different maybe the last one this is multplication and divison equations

sammixboo (sammixboo):

Well you gotta think 59.90/9 = p So 9 * p = 54.90

OpenStudy (anonymous):

ahh i see

sammixboo (sammixboo):

:)